OpenStudy (anonymous):

David knew he made a mistake when he calculated that Gilda walks 123 miles to the station. Read through David's calculations: Using d = rt, the distance is the same, but the rate and time are different. If Gilda misses the train, it means the time t needs 7 more minutes so d = 3(t + 7). If she gets to the station 5 minutes early means the time t can be 5 minutes less so d = 4(t - 5). 3(t + 7) = 4(t - 5) 3t + 21 = 4t - 20 t = 41 d = rt, so d = 3(41) = 123 Find David's mistake in his calculations. In two or more complete sentences, explain his mistake.

OpenStudy (anonymous):

Gilda walks to the train station. If she walks at the rate of 3 mph, she misses her train by 7 minutes. However, if she walks at the rate of 4 mph, she reaches the station 5 minutes before the arrival of the train. Find the distance Gilda walks to the station.

OpenStudy (anonymous):

The time Gilda needs to get to the train station while walking with rate r = 3 is equal to (t + 7) not t. Davis in his calculation using just t. So if he puts t +7 instead of t he will get the correct answer that is d = 3(41 + 7) = 3*48 = 144 mi